bdrammeh:

A 12.0 V voltage is applied to a wire with a resistance of 3.50 ohms. what is the magnetic field 0.150 m from the wire? ____x10^____T

HanAkoSolo:

@Hero

bdrammeh:

wats the answer

TheSmartOne:

We got voltage and we have resistance, and we want magnetic field oh and we're given a distance from the wire as well Now, do you know any formula to calculate magnetic field, B? Well, you probably don't. I don't remember it either LOL but I went ahead and pulled it out for you to save time \(\text{B} =\frac{\mu _{0}\text{I}}{2\pi \text{r}}\) B is our magnetic field \(\mu_{0} =4π10^{−7}~\frac{\text{T}\cdot \text{ m}}{\text{A}}\) I is current and r is the shortest distance to the wire

TheSmartOne:

We have voltage and we have resistance. Can you calculate the current? Once you do that, you can simply plug in the current that you found, the constant (\(\mu{0}\)), and r which is 0.150 m And so it'll be a simple plug and chug

bdrammeh:

hello

bdrammeh:

lk

bdrammeh:

put the answer in ___x10^___ that form

bdrammeh:

so the answer is 4x10^-6